OpenStudy (anonymous):

y' + (1/4)y = -t^2 e^t y^5

OpenStudy (amistre64):

let z = y^(-4) maybe

OpenStudy (amistre64):

\[y' + (1/4)y = -t^2 e^t y^5\] \[y'y^{-5} + (1/4)y^{-4} = -t^2 e^t \] \[z=y^{-4}~:~z' = -4y'y^{-5}\] \[-\frac14 z'+ (1/4)~z = -t^2 e^t \]

OpenStudy (amistre64):

solve for z to determine y

OpenStudy (amistre64):

to clean it up some: \[z'-z = 4t^2 e^t\]